OpenStudy (anonymous):

What is the probability of rolling a 2 or a 3 with a six sided die?

OpenStudy (anonymous):

2/6=1/3

OpenStudy (solomonzelman):

6 sides dice, not die :)

OpenStudy (solomonzelman):

numbers 2 and 3 are 2 possibilities out of 6 possibilities. Thus we get 2/6 = 1/3

OpenStudy (anonymous):

whats the difference @SolomonZelman ?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

1/6+1/6=1/3

OpenStudy (anonymous):

u didnt understand @King.Liang

OpenStudy (anonymous):

see prob of getting a no 2=1/6

OpenStudy (anonymous):

prob of getting 3=1/6